OpenStudy (anonymous):

Statistics question with given datas of gender(M&F) and level of studied degree(Bachelor's, Master's, Professional's). What is the conditional probability that you choose a woman, given that the person chosen received a professional degree? Are the events independent?

OpenStudy (unklerhaukus):

where is the data?

OpenStudy (anonymous):

#'s of female with bachelor's degree 855, with Master's 356, with professional's 44 #'s of male with bachelor's degree 631, with Master's 238, with professional's 44

OpenStudy (anonymous):

The conditional probability formula is the P(B|A) = P(BA)/P(A) where B=female and A=professional degree. Based on your data the P(BA)=44/2168 and P(A)=88/2168 Divide these two numbers and you have your answer

OpenStudy (anonymous):

So, are the events are independent? Explain please.

OpenStudy (anonymous):

No, the events are not independent. The way to check this is to see if the P(A|B)=P(A) and if the P(B|A)=P(B). Since in this problem this is not true, the events are not independent.

OpenStudy (sirm3d):



OpenStudy (anonymous):

In the second row of your table you miscalculated 631+238+44. It should be 913, not 713.

OpenStudy (sirm3d):

i stand corrected. but as to the problem, the sum 913 becomes irrelevant.